OpenStudy (rawfeeds-):

What is the slope-intercept form of the function that contains the point (8, 12) and has a slope of −2? y = x +

OpenStudy (rawfeeds-):

@Mehek14

OpenStudy (rawfeeds-):

hey

OpenStudy (rawfeeds-):

@Photon336

OpenStudy (photon336):

so yeah we know that y = m(x)+b

OpenStudy (help108):

@Photon336 i need help on a question i tagged u on plezzz after u finish this one

OpenStudy (photon336):

slope = -2 = m y = -2x+b but we need to find b we can do this by putting the point (8,12) into the equation. -2(8)+b = 12 to find b 12 = -16+b

OpenStudy (rawfeeds-):

@Photon336

OpenStudy (photon336):

@RawFeeds- solve for b and show me what you get

OpenStudy (rawfeeds-):

i need help

OpenStudy (photon336):

12 = -16+b do you know how to solve for b?

OpenStudy (rawfeeds-):

no

OpenStudy (photon336):

all you do is add 16 to both sides 12+16 = b b = 28 so it becomes this: y = -2x+28

OpenStudy (rawfeeds-):

so thats the answer

OpenStudy (photon336):

that's what they mean by slope intercept form

OpenStudy (rawfeeds-):

ok @Photon336

OpenStudy (rawfeeds-):

ok im stuck

OpenStudy (photon336):

so we have y = -2x+28 so we are done

OpenStudy (rawfeeds-):

ok